OpenStudy (anonymous):

I have to solve this to make one side look like the other. [ sin(a+B) / sin(a-B) ] = [ tan(a) + tan(B) / tan(a) - tan(B) ]

OpenStudy (anonymous):

the "a" is alpha and the "B" is beta

OpenStudy (anonymous):

Here we need to use formula \[\sin(A+B)=sinAcosB+cosAsinB\]and\[\sin(A-B)=sinAcosB-cosAsinB\]

OpenStudy (anonymous):

yes i did that but im stuck from there. i don't know what to do next

OpenStudy (noelgreco):

Now divide numerator and denominator by cos a cos b.

OpenStudy (anonymous):

why?

OpenStudy (noelgreco):

Well, first, you can do that, and second the answer you seek will pop right out. It's just algebra.

OpenStudy (anonymous):

oh ok i see. thanks so much!

OpenStudy (noelgreco):

Sure!